# 1]Terra Quantum AG, 9000 St. Gallen, Switzerland Superposed Parameterised Quantum Circuits

### v1

@@ Abstract
Quantum machine learning has shown promise for high-dimensional data analysis, yet many existing approaches rely on linear unitary operations and shared trainable parameters across outputs. These constraints limit expressivity and scalability relative to the multi-layered, non-linear architectures of classical deep networks. We introduce superposed parameterised quantum circuits to overcome these limitations. By combining flip-flop quantum random-access memory with repeat-until-success protocols, an superposed parameterised quantum circuit embeds an exponential number of parameterised sub-models in a single circuit and induces polynomial activation functions through amplitude transformations and post-selection. We provide an analytic description of the architecture, showing how multiple parameter sets are trained in parallel while non-linear amplitude transformations broaden representational power beyond conventional quantum kernels. Numerical experiments underscore these advantages: on a 1D step-function regression a two-qubit superposed parameterised quantum circuit cuts the mean-squared error by three orders of magnitude versus a parameter-matched variational baseline; on a 2D star-shaped two-dimensional classification task, introducing a quadratic activation lifts accuracy to 81.4% and reduces run-to-run variance three-fold. These results position superposed parameterised quantum circuits as a hardware-efficient route toward deeper, more versatile parameterised quantum circuits capable of learning complex decision boundaries.
@@ 2.2 Parameterised Quantum Circuits
QML seeks to exploit quantum processors to achieve performance or expressivity advantages in tasks such as classification, regression and generative modelling [50, 51, 52]. Among the various QML paradigms, parameterised quantum circuits (PQCs) — also known as variational quantum circuits or quantum neural networks — have become the tool of choice for near-term devices [4, 53, 7]. A PQC realises a trainable function by embedding both data-dependent angles and continuous parameters into a quantum circuit and extracting classical information via measurement [54, 5, 12].
@@ 6 Conclusion and Future Works
Future work can expand in several directions. First, refining RUS protocols to reduce shot overhead and post-selection failures could ease the demands on near-term quantum devices. Second, investigating partial post-selection strategies or approximate polynomial layers may further mitigate resource costs. Third, the inclusion of error mitigation or quantum error correction methods aligned with FFQRAM-based routing could bolster circuit fidelity against hardware noise [52]. Finally, adapting SPQCs to more advanced data-encoding schemes or larger problem instances — where true quantum advantage may emerge — remains an important milestone. We anticipate that continued research on multi-layered, non-linear quantum architectures will play a central role in realizing robust and expressive quantum models in practical machine learning applications.

### v2

# Superposed parameterised quantum circuits

@@ Abstract
Quantum machine learning has shown promise for high-dimensional data analysis, yet many existing approaches rely on linear unitary operations and shared trainable parameters across outputs. These constraints limit expressivity and scalability relative to the multi-layered, non-linear architectures of classical deep networks. We introduce superposed parameterised quantum circuits to overcome these limitations. By combining flip-flop quantum random-access memory with repeat-until-success protocols, a superposed parameterised quantum circuit embeds an exponential number of parameterised sub-models in a single circuit and induces polynomial activation functions through amplitude transformations and post-selection. We provide an analytic description of the architecture, showing how multiple parameter sets are trained in parallel while non-linear amplitude transformations broaden representational power beyond conventional quantum kernels. Numerical experiments underscore these advantages: on a 1D step-function regression a two-qubit superposed parameterised quantum circuit cuts the mean-squared error by three orders of magnitude versus a parameter-matched variational baseline; on a 2D star-shaped two-dimensional classification task, introducing a quadratic activation lifts accuracy to 81.4% and reduces run-to-run variance three-fold. These results position superposed parameterised quantum circuits as a hardware-efficient route toward deeper, more versatile parameterised quantum circuits capable of learning complex decision boundaries.
@@ 2.2 Parameterised Quantum Circuits
QML seeks to exploit quantum processors to achieve performance or expressivity advantages [50] in tasks such as classification, regression and generative modelling [51, 52, 53]. Among the various QML paradigms, parameterised quantum circuits (PQCs) — also known as variational quantum circuits or quantum neural networks — have become the tool of choice for near-term devices [4, 54, 7]. A PQC realises a trainable function by embedding both data-dependent angles and continuous parameters into a quantum circuit and extracting classical information via measurement [55, 5, 12].
@@ 6 Conclusion and Future Works
Future work can expand in several directions. First, refining RUS protocols to reduce shot overhead and post-selection failures could ease the demands on near-term quantum devices. Second, investigating partial post-selection strategies or approximate polynomial layers may further mitigate resource costs. Third, the inclusion of error mitigation or quantum error correction methods aligned with FFQRAM-based routing could bolster circuit fidelity against hardware noise [53, 76]. Finally, adapting SPQCs to more advanced data-encoding schemes or larger problem instances — where true quantum advantage may emerge — remains an important milestone. We anticipate that continued research on multi-layered, non-linear quantum architectures will play a central role in realizing robust and expressive quantum models in practical machine learning applications.